To cover 14 km distance , anubhav cycles at a certain average speed. If he reduces his average speed by1 km/hr , he takes 20 min more to cover the same distance. find the original speed.
Answer by stanbon(75887)

To cover 14 km distance , anubhav cycles at a certain average speed. If he reduces his average speed by1 km/hr , he takes 20 min more to cover the same distance. find the original speed.
-------------------
1st trip DATA:
distance = 14 km ; speed = x km/hr ; time = 14/x hrs
-------
2nd trip DATA:
distance = 14 km ; speed (x-1) km/hr ; time = 14/(x-1) hrs
-------
Equation:
time - time = 1/3 hr
----
14/(x-1) - 14/x = 1/3
----
3*14x - 3*14(x-1) = x(x-1)
42x - 42x - 42 = x^2 - x
---
x^2 -x + 42 = 0
Factor:
(x-7)(x+6) = 0
Positive solution:
x = 7 km/hr (original speed)
==================================
Cheers,
Stan H.
==================